ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 210, 2015

Asperó, D. and M. A. Mota A generalization of Martin’s axiom  193

Bergelson, V., M. Einsiedler and J. Tseng Simultaneous dense and nondense orbits for commuting maps

23

Berkovich, V. G. Finiteness theorems for vanishing cycles of formal schemes

147

Binyamini, G. and D. Novikov Multiplicity operators  101

Dolfi, S., D. Gluck and G. Navarro On the orders of real elements of solvable groups

1

Einsiedler, M., S. Kadyrov and A. Pohl Escape of mass and entropy for diagonal flows in real rank one situations

245

Einsiedler, M. See V. Bergelson Ferreira, V. O. and J. Z. Gonçalves Free symmetric and unitary

pairs in division rings infinite-dimensional over their centers

297

García-Melián, J. and L. Iturriaga Multiplicity of solutions for some semilinear problems involving nonlinearities with zeros

233

Gluck, D. See S. Dolfi Gonçalves, J. Z. See V. O. Ferreira Hui, C. Y., M. Larsen and A. Shalev The Waring problem for Lie

groups and Chevalley groups

81

Iturriaga, L. See J. García-Melián Kadyrov, S. See M. Einsiedler Larsen, M. See C. Y. Hui Lindsey, K. A. Counting invariant components of hyperelliptic trans-

lation surfaces

125

Loughran, D. Rational points of bounded height and the Weil restriction

47

Martin, F. A note on tropicalization in the context of Berkovich spaces

323

Mota, M. A. See D. Asperó Navarro, G. See S. Dolfi Novikov, D. See G. Binyamini Peckner, R. Uniqueness of the measure of maximal entropy for the

squarefree flow

335

Pohl, A. See M. Einsiedler Poliakovsky, A. On the Γ-limit of singular perturbation problems

with optimal profiles which are not one-dimensional. Part II: The lower bound

359

Shalev, A. See C. Y. Hui Szostakiewicz, M., M. Urbański and A. Zdunik Fine inducing

and equilibrium measures for rational functions of the Riemann sphere

399

Tseng, J. See V. Bergelson Urbański, M. See M . Szostakiewicz Winter, D. Mixing of frame flow for rank one locally symmetric

spaces and measure classification

467

Zdunik, A. See M. Szostakiewicz